OpenStudy (anonymous):

the expression x^2(2x)^6/x^5 equals cx^e the coefficient c is ? and exponent e is?

OpenStudy (campbell_st):

ok... so start by simplifying the numerator... and remember multiplying the same base... add the powers \[\frac{x^2 \times 2^6 \times x^6}{x^5} = \frac{2^6x^8}{x^5}\] when dividing the same base, subtract the powers \[\frac{2^6x^8}{x^5} = 2^6x^{~8 - 5}\] just finish simplifying it... for the answer, and the values of c and e